# Algorithmic Fairness amid Social Determinants: Reflection, Characterization, and Approach

### v1

@@ 3.1 Assumptions in Our Analyses
For clear illustration through closed-formula theoretical derivation, we incorporate certain quantitative assumptions in our theoretical analyses of different admission procedures.
@@
Let us denote the sensitive attribute as $A$, where $a\in\mathcal{A}$ denotes under-represented minority (URM) applicant group, and $a^{\prime}\in\mathcal{A}$ denotes non-URM applicant group. There are two regions where applicants reside in, rich and poor regions, with different demographic compositions from URM/non-URM groups,
$\small\begin{array}[]{rcc}&\text{poor region}&\text{rich region}\\ \text{URM applicants}&n_{a}^{(\mathrm{poor})}&n_{a}^{(\mathrm{rich})}\\ \text{Non-URM applicants}&n_{a^{\prime}}^{(\mathrm{poor})}&n_{a^{\prime}}^{(\mathrm{rich})}\end{array}\raisebox{-11.62495pt}{,}$
@@
    the definition of “underrepresented minority”, i.e., $n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})}<n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})}$.
Condition (1) specifies that URM applicants are relatively more concentrated in the less well-off region due to historical injustice [115, 107, 6]. Condition (2) holds by definition, i.e., the total number of URM applicants is smaller than that for non-URM applicants.
@@
While there can be dependence between Race and Academic Preparedness (without conditioning on Address Region) due to historical injustice [115, 107, 6], such dependence does not indicate that Race is a determinant of applicant’s Academic Preparedness. Assumption 3.2 specifies that after conditioning on applicant’s address region, applicant’s academic preparedness is irrelevant to the demographic group. In other words, Address Region encloses region-specific social determinants related to academic preparedness, for instance, the availability of educational resources in the area and the environmental impacts on applicant’s health, but Race is not an inherent determinant of applicant’s academic preparedness
Assumption 3.3 (Gamma Parameterization of Academic Preparedness Distribution).
[…]
In Assumption 3.3, the conversion of the score maps the domain of values $S_{\text{MIN}},S_{\text{MAX}}]$ (higher score $S$ is more competitive) to $0,\infty)$, where the closer to 0 the converted score $Q$, the more competitive. The flexibility of Gamma distributions allows us to use combinations of shape and scale parameters to capture properties of the region-specific academic preparedness distribution. We selected Gamma distributions for their flexibility in capturing the one-sided skew commonly observed in academic performance distributions, consistent with prior educational assessment research [18, 110].
@@
$g<n,\text{ where }n=n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})}+n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})},$
@@
Assumption 3.4 states that while the open-enrollment college can admit all applicants, the selective college uses score thresholds to distribute the limited admissions. As we shall see in Sections 3.23.4, the exact values of thresholds depend on the admission strategy, and have fairness implications in terms of the benefits and burdens experienced by individuals from different demographic groups, as well as regions with varying levels of affluence.
(a) Quota-based admission
(b) Holistic review with plus factors
(c) Top-percentage plan
3.2 Quota-Based Admissions
Aside from the fact that the quota-based admission procedure is rigid and mechanical [120], it fails to account for the role of social determinants which vary across regions and influence applicants’ academic preparedness in different ways. As a result, employing quota-based admission can further disadvantage non-URM applicants from less well-off areas, effectively introducing additional unfairness during the attempt to rectify historical racial injustice:
Theorem 3.5 (Quota-Based Admission Incurs Unfairness w.r.t. Non-URM in Poor Region).
@@ 3.2 Quota-Based Admissions
Under Assumptions 3.13.4, let us denote with $\eta_{\mathrm{quota}}\in\big{[}1,\frac{n}{n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})}}\big{]}$ the weighting coefficient over the natural proportion of URM applicants in population, such that the quota for URM admissions in the selective college is $\eta_{\mathrm{quota}}\cdot(\frac{n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})}}{n}g)$. Then, the quota-based admission strategy imposes a more competitive requirements (in terms of score threshold) for non-URM applicants from the poor region, than that for URM applicants from the rich region, unless the following condition on region-specific CDF’s is satisfied:
@@
$\small\begin{split}\max_{q\in[0,\infty)}\frac{F^{(\mathrm{rich})}(q)}{F^{(\mathrm{poor})}(q)}\geq\frac{(n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})})\eta_{\mathrm{quota}}}{(n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})})(1-\eta_{\mathrm{quota}})+(n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})})}\leavevmode\nobreak\ \raisebox{-7.74997pt}{.}\end{split}$ (1)

### v2

@@ 4.1 Assumptions in Our Analyses
Let us denote the sensitive attribute as $A$, where $a\in\mathcal{A}$ denotes under-represented minority (URM) applicant group, and $a^{\prime}\in\mathcal{A}$ denotes non-URM applicant group. There are two regions where applicants reside in, rich and poor regions, with different demographic compositions,
$\small\begin{array}[]{rcc}&\text{poor region}&\text{rich region}\\ \text{URM applicants}&n_{a}^{(\mathrm{poor})}&n_{a}^{(\mathrm{rich})}\\ \text{Non-URM applicants}&n_{a^{\prime}}^{(\mathrm{poor})}&n_{a^{\prime}}^{(\mathrm{rich})}\end{array}\raisebox{-11.62495pt}{,}\vskip-3.87498pt$
@@
    the definition of “underrepresented minority,” i.e., $n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})}<n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})}$.
Condition (1) specifies that URM applicants are relatively more concentrated in the less well-off region due to historical injustice (Sowell, 2004; Rothstein, 2017; Alexander, 2020). Condition (2) holds by definition, i.e., the total number of URM applicants is smaller than that for non-URM applicants.
@@
While there can be marginal dependence between Race and Academic Preparedness due to historical injustice (Sowell, 2004; Rothstein, 2017; Alexander, 2020), such dependence does not indicate that Race is a determinant of applicant’s Academic Preparedness. Assumption 4.2 specifies that after conditioning on applicant’s address region (and therefore, region-specific social determinants related to education), applicant’s academic preparedness is irrelevant to the demographic group. This assumption is not intended to use region as a mediating variable to explain racial discrimination in education, e.g., as in Conditional Demographic Parity frameworks (Žliobaite et al., 2011; Kamiran et al., 2013; Wachter et al., 2021). Instead, it encodes the premise that race is not an inherent factor shaping learning ability, explicitly refuting essentialist interpretations and aligning with the literature (Roberts, 2011; Kohler-Hausmann, 2018; Smedley, 2018; Kasirzadeh & Smart, 2021; Delgado & Stefancic, 2023; Doh et al., 2025).
For tractability, Assumption 4.2 simplifies the underlying data generating processes by focusing on inter-regional structural injustice and abstracting from intra-regional and within-institution racial discrimination (e.g., bias operating within the same school), which remain important mechanisms beyond the scope of our model. Even in absence of these mechanisms, we show that the interplay between sensitive-attribute-centered mitigation and progress in structural justice remains non-trivial.
Assumption 4.3 (Stochastic Dominance of Academic Preparedness Distribution).
Let $S$ denote the non-negative overall academic index score of an applicant’s academic preparedness. Further let $S_{\text{MAX}}$ and $S_{\text{MIN}}$ denote the highest and lowest possible values of the score. The rich region’s cumulative distribution function (CDF), of log-converted relative score $Q$ dominates that of the poor region:
$\small\begin{split}&\text{Let }Q\coloneqq-\log\big(\frac{S-S_{\text{MIN}}}{S_{\text{MAX}}-S_{\text{MIN}}}\big),\text{ then for }r\in\{\text{poor},\text{rich}\},\\ &\text{the CDF's satisfy }F^{(\mathrm{rich})}(q)\geq F^{(\mathrm{poor})}(q),\forall q\in[0,\infty).\end{split}$
The degree of stochastic dominance in academic preparedness between applicants from rich and poor regions captures the level of structural injustice.
@@
$\small g<n,\text{ where }n=n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})}+n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})},$
@@
4.2 Structural Injustice Implications of A Classic Mitigation Strategy
The quota-based admission is a type of affirmative-action admission strategy, originally designed to rectify historical injustice by directly setting aside admission quotas to increase the representation of URM students. Aside from the fact that the quota-based admission procedure is rigid and mechanical (Grutter v. Bollinger (2003) Supreme Court, 2003b), this sensitive-attribute-centered mitigation strategy fails to account for the role of social determinants, which vary across regions and influence academic preparedness.
Our results quantitatively recover the well-known concern that quota-based admissions reduce opportunities for non-URM students from disadvantaged regions by shrinking the pool of open seats. Moreover, we uncover a counter-intuitive paradox: this unintended burden is less likely under more severe structural injustice, and more likely as conditions improve.
Theorem 4.5 (Tension Between Structural Justice and Sensitive-Attribute-Centered Mitigation).
@@ 4.2 Structural Injustice Implications of A Classic Mitigation Strategy
Under Assumptions 4.14.4, let us denote with $\eta_{\mathrm{quota}}\in\big[1,\frac{n}{n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})}}\big]$ the weighting coefficient over the natural proportion of URM applicants in population, such that the quota for URM admissions in the selective college is $\eta_{\mathrm{quota}}\cdot(\frac{n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})}}{n}g)$. Then, the quota-based admission strategy imposes a more competitive requirements (in terms of score threshold) for non-URM applicants from the poor region, than that for URM applicants from the rich region, unless the following condition is satisfied:
@@
$\small\begin{split}\max_{q\in[0,\infty)}\frac{F^{(\mathrm{rich})}(q)}{F^{(\mathrm{poor})}(q)}\geq\frac{\eta_{\mathrm{quota}}}{1+(1-\eta_{\mathrm{quota}})\frac{n_{a}^{(\mathrm{poor})}+n_{a}^{(\mathrm{rich})}}{n_{a^{\prime}}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{rich})}}}\penalty 10000\ \raisebox{-9.68745pt}{.}\end{split}$ (1)
In simple terms, Theorem 4.5 establishes that (i) quota-based mitigation can disadvantage non-URM applicants from disadvantaged regions, and (ii) counter-intuitively, such harm is easier to avoid when structural injustice is more severe, and more aggressive sensitive-attribute-centered mitigation only makes generating new injustice more likely.